OpenStudy (anonymous):

Establish the identity: (sec theta/csc theta) + (sin theta/ cos theta)= 2 tan theta.

zepdrix (zepdrix):

Oh that's a fun one :) So do you see the identity in the second term?

zepdrix (zepdrix):

\[\Large\rm \frac{\sec \theta}{\csc \theta}+\color{red}{\frac{\sin \theta}{\cos \theta}}\]The red one should jump right out at you! :O

OpenStudy (anonymous):

Yes, I thought about going in that route, but I wasn't sure if it was the correct route to go. Thank you so much! You've helped me with so many questions today!

zepdrix (zepdrix):

Ah :) np